OpenStudy (anonymous):

how do i simplify this, assuming all variables are real numbers? ³√64(x-8)³ (the radical sign extends over the whole equation)

OpenStudy (anonymous):

\(\Large \sqrt[3] {64 \cdot (x-8)^3}=\sqrt[3] {64} \cdot \sqrt[3] {(x-8)^3}=4(x-8) \)

OpenStudy (anonymous):

Thank you. And, how did you get the mathematical symbols?

OpenStudy (anonymous):

right-click on the symbols themselves and choose "Show math as" --> "tex commands". it's like entering equations using the blue "Equation" button on the left side of this reply box.